The width of a computer monitor is 31 cm.if the perimeter is 148 cm, what is the area of the computer monitor
Answer by macston(5194)

The width of a computer monitor is 31 cm.if the perimeter is 148 cm, what is the area of the computer monitor
H=height; W=width=31cm; P=perimeter=2(H+W)=148cm
P=2(H+W) Divide each side by 2
P/2=H+W Subtract W from each side.
P/2-W=H
(148 cm/2)-31 cm=H
74 cm-31 cm=H
43 cm=H Height of monitor is 43 cm
Area=H*W=43 cm*31 cm=1333 sq cm ANSWER: The area of the monitor is 1333 sq cm.